### v1

# Convergence of patterned matrices with random walk entries

Abstract
It is well known that the Brownian motion on the real line can be obtained as a weak limit of a suitably scaled continuous-time random walk (CTRW). We investigate the convergence of certain patterned random matrices whose entries are independent CTRWs of various types. In a non-commutative probability framework, we use these high dimensional matrices to derive approximations of the free Brownian motion. Furthermore, we introduce and analyze a random time-changed version of the free Brownian motion driven by an inverse stable subordinator. An approximation of this process is obtained using a random matrix whose entries consist of continuous-time randomly stopped random walks. Moreover, it is shown that the empirical spectral distributions of such matrices have longer tails. Additionally, in a specific case, we use the explicit eigenvalue expressions of these matrices to obtain weak approximations of the standard Brownian motion and a time-changed variant of it.
AMS Subject Classification [2020]: Primary: 60B10; Secondary: 60B20, 15B52
Keywords: Patterned matrix, Wigner matrix, fractional Poisson process, semi-circle law, Brownian motion, free Brownian motion, inverse stable subordinator.
1 Introduction
Let $\{X_{k}\}_{k\geq 1}$ be independent and identically distributed (iid) random variables (rvs) with mean zero and variance one, and let $\{S_{n}:=\sum_{k=1}^{n}X_{k}\}_{n\geq 0}$ with $S_{0}=0$. It is well known that the scaled continuous-time random walk (CTRW), $\{n^{-1/2}S_{[nt]}\}_{t\geq 0}$, $n\geq 1$, where $\cdot$] is the greatest integer function, converges weakly to a (standard) Brownian motion (SBM) $\{W(t)\}_{t\geq 0}$, as $n\rightarrow\infty$. This conclusion remains true if $\{X_{k}\}$ are not necessarily identically distributed but are independent with uniformly bounded moments of all orders. The SBM is a Gaussian process with independent and stationary increments, $W(0)=0$ and for all $0\leq s\leq t<\infty$, W(t)W(s)W(t)-W(s) is a Gaussian variable with mean zero and variance tst-s.
Let $\{N(t)\}_{t\geq 0}$ be a renewal process with iid inter-event times $\{J_{k}\}_{k\geq 1}$, independent of $\{X_{k}\}_{k\geq 1}$. Then $\{S_{N(t)}\}_{t\geq 0}$ is a time-changed random walk. If $\{N(t)\}_{t\geq 0}$ is a Poisson process with rate one and compounding rvs $\{X_{k}\}_{k\geq 1}$ are standard Gaussian then the scaled CTRW, $\{n^{-1/2}S_{N(nt)}\}_{t\geq 0}$, $n\geq 1$, also converges weakly in an appropriate Skorokhod topology to an SBM (see [26]). Moreover, if $\{J_{k}\}_{k\geq 1}$ belong to the domain of attraction of a stable law with index $\alpha\in(0,1)$ and $\{X_{k}\}_{k\geq 1}$ belong to the domain of attraction of a standard Gaussian law, then the scaled CTRW, $\{n^{-\alpha/2}S_{N(nt)}\}_{t\geq 0}$, $n\geq 1$, $0<\alpha<1$, weakly converges in an appropriate Skorokhod topology to a time-changed Brownian motion (see [23, 26]), where the time changing component is an inverse stable subordinator (for definition see Section 2.4).
In non-commutative probability, the analogues of the Gaussian distribution and the Brownian motion are respectively the semi-circle distribution (also known as the free Gaussian distribution) and the (standard) free Brownian motion (FBM), say $\{B(t)\}_{t\geq 0}$ (see Section 2.2) for which the increments are freely independent and the moments of $B(t)$ coincides with the moments of the semi-circle distribution with variance $t$. At the same time, it is known that the algebraic limit variable of scaled Wigner matrices of increasing dimensions has the semi-circle distribution, and that independent Wigner matrices are almost surely (a.s.) asymptotically free.
Symmetric matrix valued random processes and their associated eigenvalue processes have been studied in great detail for many years since their formal introduction in [14]. The connection between positive semi-definite matrix valued random processes and the dynamical systems associated with their eigenvalues is discussed in [25]. For more details on the recent development for the eigenvalues of matrix-valued random processes, we refer to [27, 28, 29], and references therein. In [4], FBM is obtained as a scaling limit of Wigner matrices with independent Brownian motion entries. In the last few decades, the connection between symmetric matrix valued random processes and multivariate statistical analysis and their application have been investigated in various fields, such as financial data analysis (see [15, 16, 17]), computer vision (see [20]), and machine learning (see [34]).
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a classical probability space. Let $\mathcal{A}_{\mathbb{P}}=\cap_{1\leq p<\infty}L^{p}(\mathbb{P})$ be the algebra of rvs with all moments finite and let $\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}})$ be the algebra of $n\times n$ complex random matrices with entries from $\mathcal{A}_{\mathbb{P}}$. Then, the pair $(\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}}),\tau_{n})$ with $\tau_{n}=n^{-1}\mathbb{E}^{\mathbb{P}}\mathrm{Trace}(\cdot)$ or $\tau_{n}=n^{-1}\mathrm{Trace}(\cdot)$ forms a non-commutative probability space (see Section 2.1 for a formal definition). Let $\{A_{n}(t)\}_{t\geq 0}$ be a collection from $\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}})$. Then, it is said to be algebraically convergent to a collection $\{A(t)\}_{t\geq 0}$ in finite dimensional sense if for any $p\in\mathbb{N}$ and distinct $t_{1},\dots,t_{p}\in[0,\infty)$, the matrices $A_{n}(t_{1}),\dots,A_{n}(t_{p})$ jointly converge to variables $A(t_{1}),\dots,A(t_{p})$ (see Section 2.1.2 for details).
Motivated by the above developments, we consider some sequences of patterned random matrices whose entries evolve as independent CTRWs. Let $(S_{i,j,n})_{1\leq i,j\leq n}$, $n\geq 1$ be independent random walks whose steps are independent with mean 0, variance $1$ and all moments uniformly bounded. For $d\in\{1,2\}$, let $L_{n}:\{1,2,\dots,n\}^{2}\rightarrow\mathbb{Z}^{d}$ be link functions, with $L_{n+1}(i,j)=L_{n}(i,j)$ for all $1\leq i,j\leq n$. By abuse of notation, we denote these functions by $L$. For a suitable symmetric $L$ (that is, $L(i,j)=L(j,i)$ for all $i,j$), these yield symmetric random matrices $\{A_{n}(t)=(S_{L(i,j),[nt]})_{1\leq i,j\leq n}\}_{t\geq 0}$ in the NCP $(\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}}),\tau_{n})$. We also consider a time-changed version of this process $\{A_{n}^{\alpha}(t)=(S_{L(i,j),N^{\alpha}(t)})_{1\leq i,j\leq n}\}_{t\geq 0}$, $0<\alpha\leq 1$, where the time-changing component $\{N^{\alpha}(t)\}_{t\geq 0}$ is an independent fractional Poisson process (see Section 2.4 for the definition).
After recalling in Section 2, some preliminaries that we need, in Sections 3.1 and 3.2, we establish respectively, the marginal and finite dimensional joint algebraic convergence as defined above (that is, convergence of the expected average trace of any polynomial) of $\{n^{-1}A_{n}(t)\}_{t\geq 0}$ for suitable links $L$. These algebraic convergence results also yield the convergence of the empirical spectral distribution of symmetric polynomials in independent copies of these matrices, evaluated at different time points.
In particular, for the choice of $L(i,j)=(\max(i,j),\min(i,j))$, $\{n^{-1}A_{n}(t)\}_{t\geq 0}$ converges algebraically to a FBM $\{B(t)\}_{t\geq 0}$. The empirical spectral distributions of any symmetric polynomial in these matrices converges to the distributions determined by the moments of the corresponding (self-adjoint) polynomial of $\{B(t)\}_{t\geq 0}$. The choice of L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$, yields a Symmetric Circulant matrix. In this case, the matrices are commutative and $\{n^{-1}A_{n}(t)\}_{t\geq 0}$ converge algebraically to $\{W(t)\}_{t\geq 0}$, whose finite dimensional moments agree with those of an SBM.
In Section 3.3, with the link function L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$, we consider the empirical process, $Y_{n}(t)=n^{-1}\sum_{r=1}^{n}\lambda_{r,n}(t)\mathbb{I}(U_{n}=r)$ where $\lambda_{r,n}(t),1\leq r\leq n$ are the eigenvalues of $A_{n}(t)$ and $U_{n}$ is a uniform random variable independent of all other variables. We show that $Y_{n}(\cdot)$ converges weakly to an SBM.
In Section 4 we introduce the time-changed FBM. In Section 4.1, we consider patterned random matrices with stopped (via fractional Poisson process) random walk entries. we establish the finite dimensional algebraic convergence of the matrices $n^{-(1+\alpha)/2}A_{n}^{\alpha}(t)_{t\geq 0}$. In particular, for the choice $L(i,j)=(\max(i,j),\min(i,j))$ the limit is $\{B(L^{\alpha}(t))\}_{t\geq 0}$, where $\{L^{\alpha}(t)\}_{t\geq 0}$ is an inverse stable subordinator independent of $B$. The moments of $B(L^{\alpha}(t))$ identify a unique probability distribution. Simulations (see Figure 1) show that its density function has a longer tail compared to the semi-circle density with variance $t$, and as $\alpha\to 1$, the value of $\alpha$ increases, this density approaches the latter. In Section 4.2 we consider the link function L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$, and prove results similar to those of Section 3.3, the marginal limit in this case being $\{W(L^{\alpha}(t))\}_{t\geq 0}$, where $\{L^{\alpha}(t)\}_{t\geq 0}$ is an inverse stable subordinator independent of $W(t)$.
2 Notations and preliminaries
In this section, we fix our notation and recall known concepts and results that will be used later. Let $\mathbb{C}$, $\mathbb{Z}$ and $\mathbb{N}$ denote the sets of complex numbers, integers and positive integers, respectively. For any $\subset\mathbb{Z}, #B\#B$ will denote its cardinality. We denote constants by $K$ or $K_{1},K_{2},\dots$, often without specifying their dependence on different variables.
2.1 Non-commutative probability and (free) independence
A non-commutative probability space (NCP) is a pair $(\mathcal{A},\tau)$ with unital algebra $\mathcal{A}$ over $\mathbb{C}$ and a linear functional $\tau:\mathcal{A}\rightarrow\mathbb{C}$ such that $\tau(1_{\mathcal{A}})=1$, where $1_{\mathcal{A}}$ denotes the identity in $\mathcal{A}$. The state $\tau$ is called tracial if $\tau(ab)=\tau(ba)$ for all $a,b$ in $\mathcal{A}$. If $\mathcal{A}$ is a *-algebra, that is, there is an involution or *-operator $\mathcal{A}\rightarrow\mathcal{A}$. A state $\tau$ on $\mathcal{A}$ is called positive if $\tau(a^{*}a)\geq 0$ for all $a\in\mathcal{A}$. The pair $(\mathcal{A},\tau)$ with *-algebra $\mathcal{A}$ and positive state $\tau$ is called *-probability space.
2.1.1 Independence and free independence
Let $(\mathcal{A},\tau)$ be an NCP and $\{\mathcal{A}_{j}\}_{j\in\mathcal{J}}$ be a collection of unital sub-algebras of $\mathcal{A}$. Then, they are called independent if they commute and for all finite subsets $\mathcal{I}\subset\mathcal{J}$ and $a_{i}\in\mathcal{A}_{i}$, $i\in\mathcal{I}$, the state $\tau$ factorizes as follows: $\tau\big(\prod_{i\in\mathcal{I}}a_{i}\big)=\prod_{i\in\mathcal{I}}\tau(a_{i})$.
The sub-algebras $\{\mathcal{A}_{j}\}_{j\in\mathcal{J}}$ are free independent if for $k=1,2,\dots,n$, $n\in\mathbb{N}$ and $a_{k}\in\mathcal{A}_{i_{k}}$, $i_{k}\in\mathcal{I}$ such that $i_{k^{\prime}}\neq i_{k^{\prime}+1}$ for $k^{\prime}=1,2,\dots,n-1$ and $\tau(a_{k})=0$ for all $k$, we have $\tau(a_{1}\dots a_{n})=0$.
Elements of an algebra are called independent (resp. free independent) if the sub-algebras generated by them are independent (resp. free independent). Appropriate extensions of these concepts are available for *-probability spaces.
2.1.2 Joint distribution and asymptotic freeness
Let $(\mathcal{A},\tau)$ be an NCP. Let $\mathbb{C}\langle x_{1},\dots,x_{p}\rangle$ be the set of all polynomials in non-commutative indeterminates $x_{1},\dots,x_{p}$, and let $g$ denote a typical polynomial in these variables. For $a_{1},\dots,a_{p}\in\mathcal{A}$, $p\geq 1$, their joint distribution is a linear functional $\varphi:\mathbb{C}\langle a_{1},\dots,a_{p}\rangle\to\mathbb{C}$ given as follows:
$\varphi(g)=\tau(g(a_{1},\dots,a_{p})).$
Let $(\mathcal{A}_{n},\tau_{n})$, $n\geq 1$ be NCPs. Let $\{a_{1,n},\dots,a_{p,n}\}\subset\mathcal{A}_{n}$, with joint distribution $\varphi_{n}$. Then, $a_{1,n},\dots,a_{p,n}$ are said to jointly converge (in the algebraic sense) if $\lim_{n\longrightarrow\infty}\varphi_{n}(g)$ exists for all fixed $g\in\mathbb{C}\langle a_{1,n}\dots,a_{p,n}\rangle$. The limits can be used in a natural way to define an NCP, say $(\mathcal{A},\varphi)$ where $\mathcal{A}$ is generated by indeterminate variables $a_{1},\ldots,a_{p}$, so that $\lim_{n\longrightarrow\infty}\varphi_{n}(g)=\varphi(g)$ for all $g$. If the limiting variables $a_{1},\dots,a_{p}$ are free independent then $a_{1,n},\dots,a_{p,n}$ are called asymptotically free. Again, obvious extensions of this concept is available for *-probability spaces.
Since we will be dealing with matrices, our *-algebra of interest will be $\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}})$ consisting of all $n\times n$ complex random matrices with entries from $\mathcal{A}_{\mathbb{P}}$, where $\mathcal{A}_{\mathbb{P}}=\cap_{1\leq p<\infty}L^{p}(\mathbb{P})$ is the algebra of all rvs defined on a given probability space with all moments finite. The states on $\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}})$ will be either $\tau_{n}=n^{-1}\mathbb{E}^{\mathbb{P}}\mathrm{Trace}(\cdot)$ or $\tau_{n}=n^{-1}\mathrm{Trace}(\cdot)$.
2.2 Free Brownian motion
A free Brownian motion (FBM) is a collection $\{B(t)\}_{t\geq 0}$ of self-adjoint operators in a $W^{*}$-probability space $(\mathcal{A},\tau)$ such that (for details see [3, 4]),
(i) $B(0)=0$;
(ii) For $p\in\mathbb{N}$ and $0\leq t_{0}<t_{1}<\dots<t_{p}<\infty$, the increments $B(t_{1})-B(t_{0}),B(t_{2})-B(t_{1}),\dots,B(t_{p})-B(t_{p-1})$ are freely independent;
(iii) For any $s,t\geq 0$, B(s+t)B(s)B(s+t)-B(s) is a semi-circle variable with variance $t$. That is, its moments uniquely determine a semi-circular law with density given as follows:
$f(x,t)=\begin{cases}\displaystyle\frac{1}{2\pi t}\sqrt{4t-x^{2}},\ x\in[-2\sqrt{t},2\sqrt{t}],\\ 0,\ \text{otherwise}.\end{cases}$
2.3 Patterned random matrices and link function
Note that the notion of convergence as defined in Section 2.1.2 applies to random matrices with the choice of state being $\mathbb{E}\text{Trace}$. A related concept of convergence is that of the spectral distribution. Let $A_{n}$ be an $n\times n$ symmetric random matrix whose entries are rvs on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, and its eigenvalues are $\lambda_{1},\dots,\lambda_{n}$. The empirical spectral distribution (ESD) of $A_{n}$ is defined as follows:
$F_{A_{n}}(x)=n^{-1}\sum_{k=1}^{n}\mathbb{I}_{\{\lambda_{k}\leq x\}},\ x\in\mathbb{R},$ (2.1)
where $\mathbb{I}$ is the indicator function and $\mathbb{R}$ denotes the set of real numbers. The expected empirical spectral distribution (EESD) of $A_{n}$ is defined as $\bar{F}_{A_{n}}(x)\coloneqq\mathbb{E}F_{A_{n}}(x)$, $x\in\mathbb{R}.$ The limiting spectral distribution (LSD) of $\{A_{n}\}_{n\geq 1}$ is the weak limit of random sequence $\{F_{A_{n}}\}_{n\geq 1}$, whenever it exist. The ESD of $A_{n}$ is said to be weakly a.s. converge to a non random distribution function $F$ if $\mathbb{P}\{\omega\in\Omega:\ \text{$F_{A_{n}}(x)\rightarrow F(x)$ as $n\rightarrow\infty$}\}=1$ for all continuity points $x$ of $F$.
Wigner matrix is one of the most important random matrices, defined as $W_{n}:=((x_{ij}))_{n\times n}$ where $x_{ij}=x_{ji}$ and they are iid with mean 0 and variance $1$. The ESD of $n^{-1/2}W_{n}$ converges a.s. to the standard semi-circle law (see [33]). Independent scaled Wigner matrices also converge algebraically (when all moments are uniformly bounded) and are asymptotically free, the limits being free semi-circle variables. See [1] or [9].
A class of patterned random matrices that includes many important matrices has been studied in a unified way by [10]. We introduce the basic ideas of this approach now. For $d\in\{1,2\}$, let us consider a sequence $\{L_{n}\}_{n\geq 1}$ of functions defined as $L_{n}:\{1,2,\dots,n\}^{2}\rightarrow\mathbb{Z}^{d}$ with $L_{n+1}(i,j)=L_{n}(i,j)$. By using a common notation $L\coloneqq L_{n}$, it is called the link function. Now, by abuse of notation, $\mathbb{N}^{2}$ is a common domain of the sequence $\{L_{n}\}$.
Let $\{a_{i},\ i\geq 1\}$ and $\{a_{i,j},\ i\geq 1,\ j\geq 1\}$ be sequence and bi-sequence of rvs. For $n\geq 1$, a $n\times n$ random matrix $A_{n}$ with link function $L$ and input sequence $\{a_{i},\ i\geq 1\}$ or $\{a_{i,j},\ i\geq 1,\ j\geq 1\}$ is defined as $A_{n}\coloneqq((a_{L(i,j)}))_{1\leq i,j\leq n}$. The following are examples of link functions for some well known random matrices:
(i) $L(i,j)=(\max\{i,j\},\min\{i,j\})$: Wigner matrix;
(ii) L(i,j)=|ij|L(i,j)=|i-j|: (symmetric) Toepliz matrix;
(iii) L(i,j)=i+j2L(i,j)=i+j-2: (symmetric) Hankel matrix;
(iv) L(i,j)=i+j2modnL(i,j)=i+j-2\ $\text${mod}\ n, $1\leq i,j\leq n$: Reverse Circulant matrix;
(v) L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$: Symmetric Circulant matrix.
For several convergence results for patterned random matrices, both for the ESD and for joint algebraic convergence, see [9, 10, 12] and references therein. The following property that all the above link functions satisfy plays a crucial role:
Property B
$\Delta L\coloneqq\sup_{n}\sup_{r\in\mathbb{Z}^{d}_{+}}\sup_{1\leq i\leq n}\#\{j:1\leq j\leq n,\ L(i,j)=r\}<\infty$.
From (2.1), the $l$th moment associated to the ESD of $A_{n}$ is $\mu^{(l)}_{F_{A_{n}}}=n^{-1}\mathrm{Trace}(A_{n}^{l})$, $l\geq 1$, and the $l$th moment associated to its EESD is given by the trace-moment formula:
$\mathbb{E}\mu_{F_{A_{n}}}^{(l)}=n^{-1}\mathbb{E}\mathrm{Trace}(A_{n}^{l})=n^{-1}\sum_{1\leq j_{1},\dots,j_{l}\leq n}\mathbb{E}a_{L(j_{1},j_{2})}\dots a_{L(j_{l},j_{1})},\ l\geq 1.$
Let us consider a non random sequence $\{\mu^{(l)}\}_{l\geq 1}$ such that
(M1) For all $l\geq 1$, $\mathbb{E}\mu^{(l)}_{F_{A_{n}}}\rightarrow\mu^{(l)}$ as $n\rightarrow\infty$;
(M2) $\mathbb{V}\mathrm{ar}(\mu^{(l)}_{F_{A_{n}}})\rightarrow 0$ as $n\rightarrow\infty$.
If the sequence $\{\mu^{(l)}\}_{l\geq 1}$ satisfies Carleman’s condition, $\sum_{l=1}^{\infty}(\mu^{(2l)})^{-1/2l}=\infty$, then $\{\mu^{(l)}\}_{l\geq 1}$ uniquely determines a probability distribution, and in that case the ESD converges in probability to this distribution. The convergence can be strengthened to a.s. sense if (M2) is replaced by the following fourth moment condition:
(M4) for all $n\geq 1$, $\sum_{l=1}^{\infty}\mathbb{E}\big(\mu^{(l)}_{F_{A_{n}}}-\mathbb{E}\mu^{(l)}_{F_{A_{n}}}\big)^{4}<\infty$.
This route has been taken for patterned random matrices. The key step is to verify (M1). We outline how this is done using the trace-moment formula. For $l\geq 1$ and $n\geq 1$, a function $\pi:\{0,1,\dots,l\}\rightarrow\{1,2,\dots,n\}$ with $\pi(0)=\pi(l)$ is called a circuit of length $l\geq 1$. Thus, with $A_{\pi}=a_{L(\pi(0),\pi(1))}\dots a_{L(\pi(l-1),\pi(l))}$,
$\mathbb{E}\mu_{F_{A_{n}}}^{(l)}=n^{-1}\sum_{\pi:\,\text{$\pi$ is circuit of length $l$}}\mathbb{E}A_{\pi}.$
A circuit $\pi$ is called matched if all the $L$-values $L(\pi(j-1),\pi(j))$ are repeated. It is called pair-matched if each value $L(\pi(j-1),\pi(j))$ appears exactly twice. We group these circuits into equivalence classes: circuits $\pi_{1}$ and $\pi_{2}$ belong to the same class, denoted as $\pi_{1}\sim\pi_{2}$ if $L(\pi_{1}(j-1),\pi_{1}(j))=L(\pi_{1}(i-1),\pi_{1}(i))$ if and only if $L(\pi_{2}(j-1),\pi_{2}(j))=L(\pi_{2}(i-1),\pi_{2}(i))$ for all $1\leq i,j\leq l$. These equivalence classes can be indexed by partitions of $\{1,\dots,l\}$, which are represented by words $w$ of length $l$ in which letters appear in alphabetical order. Let $w(i)$ denote the $i$th letter of $w$ and $\Pi(w)$ be the equivalence class associated with word $w$. Then,
$\Pi(w)\coloneqq\{\pi:w(i)=w(j)\ \text{if and only if}\ L(\pi(j-1),\pi(j))=L(\pi(i-1),\pi(i))\}.$
Hence,
$\mathbb{E}\mu_{F_{A_{n}}}^{(l)}=\sum_{w:\ \text{$w$ is a word of length $l$}}n^{-1}\sum_{\pi\in\Pi(w)}\mathbb{E}A_{\pi}.$ (2.2)
A word $w$ is called matched if each letter repeats, and it is called pair-matched if each letter repeats exactly twice. Note that a pair-matched word is always of even length. Let $\mathcal{P}_{2}(2m)$ denote the set of pair-matched word of length $2m$, $m\geq 1$. It is then shown that the inner sum has a limit for each $w$ (and is 0 if $w\notin\mathcal{P}_{2}(2m)$ for any $m$). This then gives the convergence of the ESD and also implies algebraic convergence. For details, see [9].
2.3.1 Joint convergence of independent random matrices
The same ideas can be used for establishing joint algebraic convergence. Let $A_{r,n}$, $r=1,2,\dots,p$ be $n\times n$ independent random matrices with common link function $L$. The $(i,j)$th entry of $A_{r,n}$ will be denoted by $A_{r,n}(L(i,j))$. Note that a typical monomial $g$ in terms of $\{A_{r,n}\}_{1\leq r\leq p}$ has the following form: $g(\{A_{r,n}\}_{1\leq r\leq p})=A_{r_{1},n}\dots A_{r_{l},n}$, where $r_{j}\in\{1,2,\dots,p\}$ for each $j=1,2,\dots,l$. The collection $\{A_{r,n}\}_{1\leq r\leq p}$ in $(\mathcal{M}_{n}(\mathbb{C}),\tau_{n}=n^{-1}\mathbb{E}\mathrm{Trace})$ converges jointly if for every fixed monomial $g$, the following converges:
$\tau_{n}(g)=n^{-1}\sum_{\pi:\,\text{$\pi$ is a circuit of length $m$}}\mathbb{E}A_{r_{1},n}(L(\pi(0),\pi(1)))\dots A_{r_{l}n}(L(\pi(l-1),\pi(l))).$
It turns out that monomials $g$ with odd number of factors always yield the limit 0 and hence without loss we assume that the order of $g$ is $2m$ for some integer $m$. A circuit $\pi$ of length $l$ is called index-matched if the value $(r_{j},L(\pi(j-1),\pi(j)))$ repeats for each $j=1,2,\dots,l$. Note that circuits are index-matched with reference to the monomial $g$. Let $I=\{\pi:\ \pi\ \text{is a index-matched circuit}\}$. Consider a equivalence class on $I$ where for any two circuits $\pi_{1}$ and $\pi_{2}$, we have $\pi_{1}\sim\pi_{2}$ if for any $1\leq i,j\leq l$ such that $r_{i}=r_{j}$, $A_{r_{i},n}(L(\pi_{1}(i-1),\pi_{1}(i)))=A_{r_{j},n}(L(\pi_{1}(j-1),\pi_{1}(j)))$ if and only if $A_{r_{i},n}(L(\pi_{2}(i-1),\pi_{2}(i)))=A_{r_{j},n}(L(\pi_{2}(j-1),\pi_{2}(j)))$. Any such equivalence class can be indexed by a partition of $\{1,2,\dots,l\}$ which further induces an indexed word $w$ of length $l$, where the indexed letters appear in alphabetical order. An indexed word is called pair-matched if every indexed letter in it appears exactly twice. Only such letters remain relevant in the limit. Let $w_{r_{j}}(j)$ denote the $j$th letter of indexed matched word $w$ and $\Pi_{g}(w)$ denote the class of circuits corresponding to $w$ and monomial $g$. Then,
$\Pi_{g}(w)=\{\pi:\ w_{r_{i}}(i)=w_{r_{j}}(j)\ \text{iff}\ A_{r_{i},n}(L(\pi(i-1),\pi(i)))=A_{r_{j},n}(L(\pi(j-1),\pi(j)))\}.$
The set of indexed pair-matched words ($g$ is necessarily of length $2m$) is denoted by
$\mathcal{P}_{2}(2m,g)=\{\text{all indexed pair-matched word of length $2m$, $m\geq 1$}\}.$
If we drop the indices of the indexed word $w$, then it becomes a matched word. Thus, for each monomial $g$ (of order $2m$) there exist a bijection $\chi_{g}:\mathcal{P}_{2}(2m,g)\rightarrow\mathcal{P}_{2}(2m)$.
Using the results for a single matrix, this would show the joint convergence of the iid copies of the pattern matrices described above. For further details, see [6, 9, 10, 11], and references therein.
2.4 Inverse stable subordinator and fractional Poisson process
For $0<\alpha<1$, a non-negative and non-decreasing Lévy process $\{H^{\alpha}(t)\}_{t\geq 0}$ is an $\alpha$-stable subordinator if its Laplace transform is given by $\mathbb{E}e^{-uH^{\alpha}(t)}=e^{-tu^{\alpha}}$, $u>0$ (see [24]).
A first passage time process $\{L^{\alpha}(t)\}_{t\geq 0}$ defined by $L^{\alpha}(t)\coloneqq\inf\{s\geq 0:H^{\alpha}(s)\geq t\}$ is called inverse $\alpha$-stable subordinator. Also, we set $L^{1}(t)=t$. The $l$th order moment of $L^{\alpha}(t)$ is given by
$\mathbb{E}(L^{\alpha}(t))^{l}=\frac{t^{l\alpha}\Gamma(l+1)}{\Gamma(l\alpha+1)},\ l\geq 1,\ t\geq 0.$ (2.3)
The fractional Poisson process (fPP) is a generalization of the homogeneous Poisson process. Let $\{N^{\alpha}(t)\}_{t\geq 0}$, $0<\alpha\leq 1$ be an fPP with rate $\lambda>0$ in the sense of [7]. Its state probabilities are given by $\mathrm{Pr}\{N^{\alpha}(t)=n\}=(\lambda t^{\alpha})^{n}E_{\alpha,n\alpha+1}^{n+1}(-\lambda t^{\alpha})$, $n\geq 0$, where $E_{\alpha,\beta}^{\gamma}(\cdot)$, $\alpha>0$, $\beta,\gamma\in\mathbb{R}$ is the three parameter Mittag-Leffler function (see [18], p. 47).
Let $\{N(t)\}_{t\geq 0}$ be a homogeneous Poisson process with parameter $\lambda>0$, independent of $\{L^{\alpha}(t)\}_{t\geq 0}$. Then, in view of [22], an fPP satisfies the following time-changed relationship:
$N^{\alpha}(t)\overset{d}{=}N(L^{\alpha}(t)),\ 0<\alpha\leq 1,\ t\geq 0.$ (2.4)
For more details on different characterizations of the fPP, we refer the reader to [19, 21, 22], and references therein.
3 Patterned random matrices with CTRW entries
Motivated by all the discussions so far, let $\{X_{i,j,k},i\geq 1,\,j\geq 1,\,k\geq 1\}$ be a triple index sequence of real valued rvs. Let us consider a collection of CTRWs $\{\{S_{i,j,n}(t)\}_{t\geq 0}\}_{1\leq i,j\leq n}$, $n\in\mathbb{N}$ defined as follows:
$S_{i,j}^{(n)}(t)\coloneqq\sum_{k=1}^{[nt]}X_{i,j,k},\ 1\leq i,j\leq n,\ n\in\mathbb{N}.$ (3.1)
Now we focus on certain large dimensional patterned random matrices whose entries are from (3.1), via an associated link function $L$, taking values in either $\mathbb{N}$ or $\mathbb{N}^{2}$. When $L$ takes values in $\mathbb{N}$, by a slight abuse of notation, we regard the triple indexed sequence $\{X_{i,j,k},\ i\geq 1,\,j\geq 1,\,k\geq 1\}$ as a bi-indexed sequence $\{X_{i,k},\ i\geq 1,\,k\geq 1\}$. Let us consider the sequence $\{\{A_{n}(t)\}_{t\geq 0}\}_{n\geq 1}$ of matrix valued random processes with link function $L$ defined as follows:
$A_{n}(t)=(S_{L(i,j)}^{(n)}(t))_{1\leq i,j\leq n},\ n\geq 1,\ t\geq 0.$ (3.2)
$A_{n}(0)$ is taken to be the $n\times n$ matrix of zeroes. Consider the scaled matrix $n^{-1}A_{n}(t)$, with eigenvalues are $\lambda_{1}(t),\dots,\lambda_{n}(t)$. Its ESD is given by $F_{n^{-1}A_{n}(t)}(x)\coloneqq n^{-1}\sum_{r=1}^{n}\mathbb{I}(\lambda_{r}(t)\leq x),\ x\in\mathbb{R}$, $t>0$. From (2.2), the $l$th moment of its EESD is
$\frac{\mathbb{E}\mathrm{Trace}(A_{n}(t))^{l}}{n^{1+l}}=\sum_{w:\ \text{$w$ is word of length $l$}}\frac{1}{n^{1+l}}\sum_{w\in\Pi(w)}\mathbb{E}S_{\pi}^{(n)}(t),\ l\geq 1,\ t\geq 0,$ (3.3)
where $S_{\pi}^{(n)}(t)=S_{L(\pi(0),\pi(1))}^{(n)}(t)\dots S_{L(\pi(l-1),\pi(l))}^{(n)}(t)$.
The following assumption will be needed:
Assumption I Let $\{X_{i,j,k},\ i\geq 1,\ j\geq 1,\ k\geq 1\}$ be a triple indexed sequence of independent rvs with zero mean and unit variance such that $\sup_{i,j,k}\mathbb{E}|X_{i,j,k}|^{l}<K_{l}<\infty$ for all $l\geq 3$.
3.1 Marginal convergence
Using Assumption I, the random variables $n^{-1/2}\sum_{k=1}^{n}X_{ijk}$ are independent and have uniformly bounded moments of all orders, a crucial assumption in Theorem 7.5.1 of [9]. The following two theorems follow from that theorem, and hence we omit their proofs.
Theorem 3.1.
Let $\{X_{i,j,k},\ 1\leq i\leq j\leq n,\ k\geq 1\}$, $n\in\mathbb{N}$ be a triple indexed sequence that satisfies Assumption I. Let $A_{n}(t)$ be the random matrix as defined in (3.2) with symmetric link function $L$ that satisfies Property B. Assume that the following limits exist:
$\lim_{n\rightarrow\infty}\frac{\#\Pi(w)}{n^{1+m}}\ \text{for all $w\in\mathcal{P}_{2}(2m)$, $m\geq 1$},$ (3.4)
where $\Pi(w)$ denotes the cardinality of the equivalence class $\Pi(w)$. Then, the limiting values of moments defined in (3.3) are given by
$\mu^{(l)}(t)\coloneqq\lim_{n\rightarrow\infty}\frac{\mathbb{E}\mathrm{Trace}(A_{n}(t))^{l}}{n^{1+l}}=\begin{cases}0,\ l=2m-1,\ m\geq 1,\vskip 5.69046pt\\ \displaystyle t^{m}\sum_{w\in\mathcal{P}_{2}(2m)}\lim_{n\rightarrow\infty}\frac{\#\Pi(w)}{n^{1+m}},\ l=2m,\ m\geq 1.\end{cases}$ (3.5)
Thus, the EESD of $n^{-1}A_{n}(t)$ weakly converges to the unique probability law whose $l$th moment equals $\mu^{(l)}(t)$ for each $l\geq 1$. As a consequence, the self-adjoint element $n^{-1}A_{n}(t)$ of the NCP $(\mathcal{M}_{n}(\mathbb{R}),\tau_{n}=n^{-1}\mathbb{E}\mathrm{Trace})$ converges to an element $A(t)$ in *-algebra $\mathcal{A}(t)$ generated by $A(t)$ with state $\tau$ defined as $\tau(A^{l}(t))=\mu^{(l)}(t)$.
Theorem 3.2.
For $t\geq 0$, let $A_{n}(t)$ be as in Theorem 3.1. Then,
$\mathbb{E}\big(n^{-1}\mathrm{Trace}(n^{-1}A_{n}(t))^{l}-\mathbb{E}n^{-1}\mathrm{Trace}(n^{-1}A_{n}(t))^{l}\big)^{4}=t^{2l}O(n^{-2}),\ l\geq 1,\ t\geq 0.$
Thus, (M4) holds and the ESD of $n^{-1}A_{n}(t)$ converges weakly a.s. surely to the unique probability law with moments as in (3.5).
Remark 3.1.
Let $A_{n}^{(1)},A_{n}^{(2)},\dots,A_{n}^{(n)}$ be independent $n\times n$ random matrices with input sequences $\{X_{i,j,1}\},\{X_{i,j,2}\},\dots,\{X_{i,j,n}\}$, respectively, each satisfying Assumption I, and a common symmetric link function $L$, satisfying Property B. If the limits (3.4) exist, then for each $j=1,\dots,n$, the EESD of $B_{n}^{(j)}=n^{-1/2}A_{n}^{(j)}$ weakly converges to a probability law whose $l$th moment equals $\mu^{(l)}(1)$ for every $l\geq 1$, defined in (3.5). Moreover, this convergence holds true for their ESD a.s. (see [9]).
From Theorem 3.1 and Theorem 3.2 for $t=1$, it follows that the ESD of $n^{-1/2}(B_{n}^{(1)}+\dots+B_{n}^{(n)})$ also converges weakly a.s. to the probability law whose $l$th moment is given by $\mu^{(l)}(1)$ for all $l\geq 1$. Thus, $B_{n}^{(1)},\dots,B_{n}^{(n)}$ and $n^{-1}(A_{n}^{(1)}+\dots+A_{n}^{(n)})$ have same LSD.
We now collect the LSD of some specific pattern random matrices with CTRW entries in the following corollary, In view of Theorem 3.1, its proof follows from the results of [10].
Corollary 3.1.
Let $\{X_{i,j,k},\ 1\leq i\leq j\leq n,\ k\geq 1\}$, $n\in\mathbb{N}$ be a triple indexed sequence that satisfies Assumption I. Let $\{\{A_{n}(t)\}_{t\geq 0}\}_{n\geq 1}$ be a sequence of matrix valued random processes as defined in (3.2) with link function $L$. Then,
(i) for $L$ defined by $L(i,j)=(\max\{i,j\},\min\{i,j\})$, the EESD of $n^{-1}A_{n}(t)$ converges to the semi-circular law with zero mean and variance $t$;
(ii) If L(i,j)=|ij|L(i,j)=|i-j| then the EESD of $n^{-1}A_{n}(t)$ converges to a probability law whose corresponding $(2m)$th moment is a sum of volumes of polyhedra in some $(k+1)$-dimensional hypercube;
(iii) If L(i,j)=i+j2L(i,j)=i+j-2 then the EESD of $n^{-1}A_{n}(t)$ converges to a probability law, symmetric about zero, whose corresponding $2m$th moment is a sum of volumes of polyhedra in some $(k+1)$-dimensional hypercube;
(iv) If L(i,j)=i+j2modnL(i,j)=i+j-2\ $\text${mod}\ n, $1\leq i,j\leq n$ then the EESD of $n^{-1}A_{n}(t)$ converges to a symmetric Rayleigh law whose $(2m)$th moment is given by $t^{m}m!$, $m\geq 1$;
(v) If L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$ then the EESD of $n^{-1}A_{n}(t)$ converges to the Gaussian law with mean zero and variance $t$.
3.2 Joint convergence
Now we study the finite dimensional algebraic convergence of the matrix valued process $\{n^{-1/2}A_{n}(t)\}_{t\geq 0}$. It is enough to consider the joint convergence of its increments. For $0\leq s<t<\infty$, an increment of $\{A_{n}(t)\}_{t\geq 0}$ is given by
$A_{n}(t)-A_{n}(s)\coloneqq\big(S_{i,j}^{(n)}(t)-S_{i,j}^{(n)}(s)\big)_{1\leq i,j\leq n},\ n\geq 1.$
As $\{S_{i,j}^{(n)}(t)\}_{t\geq 0}$’s have independent increments, the process $\{A_{n}(t)\}_{t\geq 0}$ also has independent increments.
For $p\in\mathbb{N}$, let $\mathbb{T}_{p}=\{t_{0},t_{1},\dots,t_{p}\}\subset[0,\infty)$ with $0\leq t_{0}<t_{1}<\dots<t_{p}$. Let $\{A(t)\}_{t\geq 0}$ be a process associated to an NCP $(\mathcal{A},\tau)$, and let
$\mathbb{T}_{p}(A(t))\coloneqq\{A_{1},\dots,A_{p}\}$ (3.6)
be the collection of increments of $\{A(t)\}_{t\geq 0}$ over $\mathbb{T}_{p}$, where $A_{r}=A(t_{r})-A(t_{r-1})$ for each $r=1,2,\dots,p$.
Theorem 3.3.
Let $\{A_{n}(t)\}_{t\geq 0}$ be as in Theorem 3.1, and let $\mathbb{T}_{p}(n^{-1}A_{n}(t))$ be the collection of increments of $\{n^{-1}A_{n}(t)\}_{t\geq 0}$ over $\mathbb{T}_{p}$ for $p\in\mathbb{N}$, as defined in (3.6). Let $g$ be a monomial given by $g(\mathbb{T}_{p}(n^{-1}A_{n}(t)))\coloneqq A_{r_{1},n}\dots A_{r_{l},n}$, where $A_{r_{j},n}\in\mathbb{T}_{p}(n^{-1}A_{n}(t))$ for each $j=1,2,\dots,l\geq 1$. Also, let us assume that the following limits exist:
$\lim_{n\rightarrow\infty}\frac{\#\Pi(\chi_{g}(w))}{n^{1+m}}\ \text{for all $w\in\mathcal{P}_{2}(2m,g)$},\ m\geq 1,$ (3.7)
where $\mathcal{P}_{2}(l,g)$ and $\chi_{g}$ are as defined in Section 2.3.1.
(i) Then,
$\lim_{n\rightarrow\infty}\frac{\mathbb{E}\mathrm{Trace}(A_{r_{1},n}\dots A_{r_{l},n})}{n}=\begin{cases}0,\ l=2m-1,\\ \displaystyle\prod_{\begin{subarray}{c}j=1\\ r_{j}\neq r_{i},\,j\neq i\end{subarray}}^{m}(t_{r_{j}}-t_{r_{j}-1})\sum_{w\in\mathcal{P}_{2}(2m,g)}\lim_{n\rightarrow\infty}\frac{\#\Pi(\chi_{g}(w))}{n^{1+m}},\ l=2m\end{cases}$ (3.8)
for all $m\geq 1$. Consequently, we can fix indeterminates $A(t)$ for every $t\geq 0$ and define an algebra $\tilde{\mathcal{A}}\coloneqq\mathrm{alg}\{\cup_{p\in\mathbb{N}}\cup_{\mathbb{T}_{p}}\mathcal{A}(\mathbb{T}_{p})\}$, generated by the union $\cup_{p\in\mathbb{N}}\cup_{\mathbb{T}_{p}}\mathcal{A}(\mathbb{T}_{p})$, where $\mathcal{A}(\mathbb{T}_{p})$ is a polynomial algebra over $\mathbb{C}$, generated by the set $\{A(t):t\in\mathbb{T}_{p}\}$, and define the state $\tilde{\tau}$ on it as follows:
$\tilde{\tau}(g(\{A(t):t\in\mathbb{T}_{l}\}))\coloneqq\lim_{n\rightarrow\infty}\tau_{n}(g(\{n^{-1}A_{n}(t):t\in\mathbb{T}_{l}\})),$ (3.9)
[…]
(ii) If $L(i,j)=(\max\{i,j\},\min\{i,j\})$ then (3.7) holds and the collection $\{A(t)\}_{t\geq 0}$ in $\tilde{\mathcal{A}}$ is a FBM. Similarly, if L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$ then the joint moments of $\{A(t)\}_{t\geq 0}$ agree with those of a SBM.
Proof.
(i) Note that the elements of $\mathbb{T}_{p}(n^{-1}A_{n}(t))$ are mutually independent. For a circuit $\pi$ of length $l$, let
$S_{\pi}^{(n)}(\mathbb{T}_{p})=\mathbb{E}(S_{L(\pi(0),\pi(1))}^{(n)}(t_{r_{1}})-S_{L(\pi(0),\pi(1))}^{(n)}(t_{r_{1}-1}))\dots(S_{L(\pi(l-1),\pi(l))}^{(n)}(t_{r_{l}})-S_{L(\pi(l-1),\pi(l))}^{(n)}(t_{r_{l}-1})).$
Then,
$n^{-1}\mathbb{E}\mathrm{Trace}(A_{r_{1},n}\dots A_{r_{l},n})=\frac{1}{n^{1+l}}\sum_{\pi:\,\pi\,\text{is circuit of length $l$}}S_{\pi}^{(n)}(\mathbb{T}_{p}).$ (3.10)
Note that if any index $r_{i}$, $i=1,\dots,l$ appears only once or circuit $\pi$ is non matched then by independent increments property and zero mean of $\{S_{i,j}^{(n)}(t)\}_{t\geq 0}$, we get $S_{\pi}^{(n)}(\mathbb{T}_{p})=0$. Thus, $S_{\pi}^{(n)}(\mathbb{T}_{p})=0$ for all non indexed matched circuits $\pi$. Let us write $M_{L(i,j)}^{(n)}(\mathbb{T}_{p})=\sum_{l^{\prime}=1}^{l}\big|S_{L(i,j)}^{(n)}(t_{r_{l^{\prime}}})-S_{L(i,j)}^{(n)}(t_{r_{l^{\prime}}-1})\big|$, $1\leq i,j\leq n$. Then,
$|S_{\pi}^{(n)}(\mathbb{T}_{p})|\leq\mathbb{E}M_{L(\pi(0),\pi(1))}^{(n)}(\mathbb{T}_{p})\dots M_{L(\pi(l-1),\pi(l))}^{(n)}(\mathbb{T}_{p}).$
For $h\geq 1$ and $s<t$, we have $\mathbb{E}|S_{L(i,j)}^{(n)}(t)-S_{L(i,j)}^{(n)}(s)|^{h}\leq K_{h}([nt]-[ns])^{[h/2]}$, which implies $\mathbb{E}(M_{L(i,j)}^{(n)}(\mathbb{T}_{p}))^{h}\leq\prod_{l^{\prime}=1}^{l}K_{h_{l^{\prime}}}([nt_{r_{l^{\prime}}}]-[nt_{r_{l^{\prime}}-1}])^{[h_{l^{\prime}}/2]}$, where $h_{1}+\dots+h_{l}=h$. Thus, for any non pair matched circuit $\pi$ of length $l$, following the arguments of the proof for Theorem 3.1, we get $|S_{\pi}^{(n)}(\mathbb{T}_{p})|\leq K_{l}\prod_{l^{\prime\prime}=1}^{l}([nt_{r_{l^{\prime}}}]-[nt_{r_{l^{\prime}}-1}])^{[h_{l^{\prime\prime}}^{\prime}/2]}$, where $h_{1}^{\prime}+\dots+h_{l}^{\prime}=l$. Now, on using $\Pi(w)\leq K_{l}n^{(l+1)/2}$ for any non pair-matched word $w$ of length $l$ (see [9], p. 120), we conclude that the contribution of a non indexed pair-matched circuit in (3.10) is zero as $n\rightarrow\infty$. Hence,
$\lim_{n\rightarrow\infty}n^{-1}\mathbb{E}\mathrm{Trace}(A_{r_{1},n}\dots A_{r_{2m},n})=\sum_{w\in\mathcal{P}_{2}(2m,g)}\lim_{n\rightarrow\infty}\frac{\#\Pi(\chi_{g}(w))}{n^{1+2m}}S_{\pi_{g,w}}^{(n)},\ m\geq 1,$
where $\pi_{g,w}\in\Pi(\chi_{g}(w))$ is a fixed indexed pair-matched circuit of length $2m$. Moreover, we have
$S_{\pi_{g,w}}^{(n)}(\mathbb{T}_{p})$ $=\mathbb{E}(S_{j_{1},j_{2}}^{(n)}(t_{r_{1}})-S_{j_{1},j_{2}}^{(n)}(t_{r_{1}-1}))^{2}\dots\mathbb{E}(S_{j_{m},j_{1}}^{(n)}(t_{r_{m}})-S_{j_{m},j_{1}}^{(n)}(t_{r_{m}-1}))^{2}$
$=([nt_{r_{1}}]-[nt_{r_{1}-1}])\dots([nt_{r_{m}}]-[nt_{r_{m}-1}]),\ w\in\mathcal{P}_{2}(2m,g),\ m\geq 1,$
where $r_{j}\neq r_{i}$ whenever $i\neq j$. So,
$\lim_{n\rightarrow\infty}\frac{\#\Pi(\chi_{g}(w))}{n^{1+2m}}S_{\pi_{g,w}}^{(n)}(\mathbb{T}_{p})=\lim_{n\rightarrow\infty}\frac{\#\Pi(\chi_{g}(w))}{n^{1+m}}(t_{r_{1}}-t_{r_{1}-1})\dots(t_{r_{m}}-t_{r_{m}-1}),\ w\in\mathcal{P}_{2}(2m,g).$
This completes the proof of (i).
(ii) Clearly $A(0)=0$. For $0\leq s<t<\infty$ and $A(s)$ and $A(t)$ in $\tilde{\mathcal{A}}$, from (3.8), we get
$\lim_{n\rightarrow\infty}\frac{\mathbb{E}\mathrm{Trace}(A_{n}(t)-A_{n}(s))^{l}}{n^{1+l}}=\begin{cases}0,\ l=2m-1,\ m\geq 1,\\ \displaystyle(t-s)^{m}\sum_{w\in\mathcal{P}_{2}(2m)}\lim_{n\rightarrow\infty}\frac{\#\Pi(w)}{n^{1+m}},\ l=2m,\ m\geq 1.\end{cases}$
Thus, from (3.5), we get $\tilde{\tau}((A(t)-A(s))^{l})=\tilde{\tau}((A(t-s))^{l})$ for all $l\geq 1$. If $L(i,j)=(\max\{i,j\},\min\{i,j\})$ then from (i) of Corollary 3.1, it follows that for all $s<t$, A(t)A(s)A(t)-A(s) is a semi-circle element with mean zero and variance tst-s. Moreover, in view of part (i), the variables in $\mathbb{T}_{p}(A(t))$, defined in (3.6), are freely independent using Theorem 9.2.1 of [9].
For the Symmetric circulant, similar arguments work. Finite dimensional convergence to moments of an SBM follows from Corollary 3.1 (v) and Theorem 9.5.1 of [9]. This completes the proof. ∎
Remark 3.2.
In [8], a general condition is obtained on the link function of random matrices that are not necessarily Wigner matrices, under which the associated LSD has the semi-circular density. Let $A_{n}(t)$ be the random matrix with link function $L$, as defined in Theorem 3.1. Define
$\beta_{n}=\max_{r}\#\{(i,j):L(i,j)=r,\ 1\leq i,j\leq n\},\ n\geq 1.$
Let us consider the following matched set between the columns $i$ and $j$ at stage $n$:
$M_{i,j}^{(n)}=\{k:1\leq k\leq n,L(k,i)=L(k,j)\},$
If $M^{*}=\sup_{n\geq 1}\sup_{1\leq i,j\leq n}\#M_{i,j}^{(n)}<\infty$, and $\beta_{n}=o(n)$, then in view of Theorem 2 of [8], the LSD of $A_{n}(t)$ has semi-circle density with variance $t>0$.
3.3 Symmetric circulant matrix and SBM
Now we consider $A_{n}(t)$’s with the link function L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$, and suppose Assumption I holds. Then $F_{n^{-1}A_{n}(t)},t\geq 0$, the ESD of $A_{n}(t),t\geq 0$ is a random process. We already know that for each $t>0$, the ESD $F_{n^{-1}A_{n}(t)}$ converges weakly a.s. to the Gaussian law with mean zero and variance $t$.
The eigenvalues $\lambda_{1,n}(t),\dots,\lambda_{n,n}(t)$ of the Symmetric Circulant matrix $n^{-1}A_{n}(t)$ are given as follows (see [13], Section 4.2):
$\lambda_{r,n}(t)=\frac{S_{0}^{(n)}}{n}+\frac{2}{n}\sum_{j=1}^{[n/2]}S_{j}^{(n)}(t)\cos{\bigg(\frac{2\pi rj}{n}\bigg)},\ t>0,\ r=1,\dots,n,$
where $S_{j}^{(n)}(t)=\sum_{k=1}^{[nt]}X_{j,k}$, as defined in (3.1).
Let $U_{n}$ be a discrete uniform rv on $\{1,2,\dots,n\}$ that is independent of $\{\lambda_{r,n}(t)\}_{t\geq 0}$ for all $n\geq 1$ and $1\leq r\leq n$. Let us now define a sequence of random processes $\{Y_{n}(t)\}_{t\geq 0}$, $n\geq 1$ as follows:
$Y_{n}(t)\coloneqq\lambda_{U_{n},n}(t)=n^{-1}\sum_{r=1}^{n}\lambda_{r,n}\mathbb{I}(U_{n}=r),\ \text{for all}\ t\geq 0.$ (3.11)
In the next result, we establish the functional convergence of $\{Y_{n}(t)\}_{t\geq 0}$, as $n\rightarrow\infty$.
Theorem 3.4.
Let $\{W(t)\}_{t\geq 0}$ be an SBM, and let $\{Y_{n}(t)\}_{t\geq 0}$ be the random process, as defined in (3.11). Then, $\{Y_{n}(t)\}_{t\in[0,1]}$ converges weakly to $\{W(t)\}_{t\in[0,1]}$, as $n\rightarrow\infty$ in the Skorokhod space $D[0,1]$.
Proof.
First, we show that the process $\{Y_{n}(t)\}_{t\geq 0}$ convergences to $\{W(t)\}_{t\geq 0}$ in finite dimensional sense, that is, $(Y_{n}(t):t\in\mathbb{T}_{p})\overset{d}{\rightarrow}(W(t):t\in\mathbb{T}_{p})$ for all $p\in\mathbb{N}$, where $\overset{d}{\rightarrow}$ denotes the convergence in distribution. For $m_{1},\dots,m_{p}\in\mathbb{N}$, we have
$\mathbb{E}\big[Y_{n}(t_{1})^{m_{1}}\dots Y_{n}(t_{p})^{m_{p}}\big]$ $=\frac{1}{n}\sum_{r=1}^{n}\mathbb{E}\lambda_{r,n}(t_{1})^{m_{1}}\dots\lambda_{r,n}(t_{p})^{m_{p}}$
$=\frac{1}{n}\mathbb{E}\mathrm{Trace}(A_{n}(t_{1})^{m_{1}}\dots A_{n}(t_{p})^{m_{p}}),$
where we have used the fact that symmetric circulant matrices commute and are simultaneously diagonalizable using the discrete Fourier matrix. Hence, the finite dimensional convergence follows from Theorem 3.3 (note that the limit moments are Gaussian moments and hence they determine the finite dimensional distributions uniquely).
Now, we need to show that the sequence $\{Y_{n}(t)\}_{t\geq 0}$, $n\geq 1$ is tight. For $0\leq s\leq t<\infty$, we have
$|\lambda_{r,n}(t)-\lambda_{r,n}(s)|=\frac{2}{n}\bigg|\sum_{j=1}^{[n/2]}\cos\bigg(\frac{2\pi rj}{n}\bigg)(S_{j}^{(n)}(t)-S_{j}^{(n)}(s))\bigg|.$
Then, for $0\leq t_{1}\leq t_{2}\leq t_{3}<\infty$, using (3.11) and the independent increments property of $\{S_{j}^{(n)}(t)\}_{t\geq 0}$, we get
$\mathbb{E}|Y_{n}(t_{3})-Y_{n}(t_{2})|^{2}$ $|Y_{n}(t_{2})-Y_{n}(t_{1})|^{2}$
$=\frac{1}{n}\sum_{r=1}^{n}\mathbb{E}|\lambda_{r,n}(t_{3})-\lambda_{r,n}(t_{2})|^{2}|\lambda_{r,n}(t_{2})-\lambda_{r,n}(t_{1})|^{2}$
$\leq\frac{4}{n^{5}}\sum_{r=1}^{n}\sum_{1\leq j_{1},j_{2}\leq[n/2]}\mathbb{E}\big|(S_{j_{1}}^{(n)}(t_{3})-S_{j_{1}}^{(n)}(t_{2}))\big|\big|(S_{j_{2}}^{(n)}(t_{3})-S_{j_{2}}^{(n)}(t_{2}))\big|$
$\ \ \cdot\sum_{1\leq j_{3},j_{4}\leq[n/2]}\mathbb{E}\big|(S_{j_{3}}^{(n)}(t_{2})-S_{j_{3}}^{(n)}(t_{1}))\big|\big|(S_{j_{4}}^{(n)}(t_{2})-S_{j_{4}}^{(n)}(t_{1}))\big|$
$\leq\frac{4}{n^{4}}\sum_{j_{1}=1}^{[n/2]}\mathbb{E}\big|(S_{j_{1}}^{(n)}(t_{3})-S_{j_{1}}^{(n)}(t_{2}))\big|^{2}\sum_{j_{3}=1}^{[n/2]}\mathbb{E}\big|(S_{j_{3}}^{(n)}(t_{2})-S_{j_{3}}^{(n)}(t_{1}))\big|^{2}$
$\leq\frac{4}{n^{4}}[n/2]([nt_{3}]-[nt_{2}])[n/2]([nt_{2}]-[nt_{1}])\leq K(t_{3}-t_{1})^{2},$
where $K$ is a constant, independent of the choice of $n$. Thus, the tightness of $\{Y_{n}(t)\}_{t\geq 0}$ follows from Theorem 13.5 and Eq. (13.14) of [5]. This completes the proof. ∎
4 Time-changed free Brownian motion
A Brownian motion time-changed by an independent inverse stable subordinator is a key mathematical model for analyzing anomalous diffusion in statistical mechanics. It is used for the modeling of motion of a particle experiencing a trapping effect. Here, we define a random time-changed FBM driven by an inverse stable subordinator. Let $\{B(t)\}_{t\geq 0}$ be an FBM on the *-probability space $(\mathcal{A},\tau)$. Let $L^{\alpha}(t),t\geq 0$ be an inverse stable subordinator on $(\Omega,\mathcal{F},\mathbb{P})$. From [31], all moments of $L^{\alpha}(t),t\geq 0$ are finite, Hence $\mathcal{A}_{\infty}$, the set of all polynomials in $L^{\alpha}(t),t\geq 0$ is an NCP with its state as expectation operator $\mathbb{E}$. Let us consider a collection $\{B^{\alpha}(t)\}_{t\geq 0}$ of variables of the tensor product space $(\hat{\mathcal{A}}:=\mathcal{A}\otimes\mathcal{A}_{\infty},\hat{\tau}:=\mathbb{E}\tau)$ defined as follows:
$B^{\alpha}(t)\coloneqq$ $B(L^{\alpha}(t)),\ 0<\alpha\leq 1,$ (4.1)
$\hat{\tau}(B^{\alpha}(t_{1})\dots B^{\alpha}(t_{l}))\coloneqq$ $\mathbb{E}\tau(B^{\alpha}(t_{1})\dots B^{\alpha}(t_{l})).$
$\hat{\tau}(B^{\alpha}(t_{1})\dots B^{\alpha}(t_{l})):=$ $\int_{\mathbb{R}^{l}_{+}}\tau(B(x_{1})\dots B(x_{l}))\mathrm{Pr}\{L^{\alpha}(t_{1})\in\mathrm{d}x_{1},\dots,L^{\alpha}(t_{l})\in\mathrm{d}x_{l}\}$
for all $l\geq 1$ and $t_{1},\dots,t_{l}\geq 0$. The collection $\{B^{\alpha}(t)\}_{t\geq 0}$ will be called a time-changed FBM. For $\alpha=1$, it reduces to the FBM.
From [31], it follows that all the finite dimensional joint moments of the inverse stable subordinator exist. Therefore, the linear functional $\hat{\tau}$ is well defined. In particular, for $l\geq 1$, we have $\hat{\tau}((B^{\alpha}(t))^{l})=0$ for all $t\geq 0$ whenever $l$ is odd and
$\hat{\tau}((B^{\alpha}(t))^{2m})=C_{m}\mathbb{E}(L^{\alpha}(t))^{m}=C_{m}\frac{t^{m\alpha}\Gamma(m+1)}{\Gamma(m\alpha+1)},\ m\geq 1,\ t\geq 0,$ (4.2)
where $C_{m}=(2m)!/(m+1)!m!$ is the $m$th Catalan number.
The following result identifies the unique distribution whose odd moments are zero and even moments are given by (4.2).
Proposition 4.1.
Let $Z(t)$ be a semi-circle rv with mean zero and variance $t\geq 0$ that is independent of an inverse $\alpha$-stable subordinator $\{L^{\alpha}(t)\}_{t\geq 0}$. Then,
$\mathbb{E}e^{uZ(L^{\alpha}(t))}=(u^{2}t^{\alpha})^{-1}\bigg(W_{\alpha,1-\alpha}(u^{2}t^{\alpha})-\frac{1}{\Gamma(1-\alpha)}\bigg),\ u\in\mathbb{R},\ t>0,$ (4.3)
where $W_{\nu,\mu}$, $\nu>-1$ is the Wright function (for definition see [18], Eq. (1.11.1)). As a consequence, $\{B^{\alpha}(t)\}_{t\geq 0}$ is a free process with $B^{\alpha}(0)=0$ and for each $t>0$, the spectral distribution of $B^{\alpha}(t)$ has a valid probability density with support $\mathbb{R}$, whose associated moment generating function (mgf) is given by (4.3).
Proof.
The mgf of $Z(t)$ is given by $\mathbb{E}e^{uZ(t)}=I_{1}(2\sqrt{t}u)/\sqrt{t}u$, $u\in\mathbb{R}$, $t>0$, where $I_{\nu}$, $\nu>-1$ is the modified Bessel function of first kind (for definition see [18], Eq. (1.7.16)). So,
$\mathbb{E}e^{uZ(L^{\alpha}(t))}$ $=\int_{0}^{\infty}\mathbb{E}e^{uZ(x)}\mathrm{Pr}\{L^{\alpha}(t)\in\mathrm{d}x\}$
$=\int_{0}^{\infty}\sum_{r=0}^{\infty}\frac{(\sqrt{x}u)^{2r+1}}{(r+1)!r!\sqrt{x}u}\mathrm{Pr}\{L^{\alpha}(t)\in\mathrm{d}x\}$
$=\sum_{r=0}^{\infty}\frac{u^{2r}}{(r+1)!r!}\int_{0}^{\infty}x^{r}\mathrm{Pr}\{L^{\alpha}(t)\in\mathrm{d}x\}$
$=\sum_{r=0}^{\infty}\frac{u^{2r}\Gamma(r+1)t^{\alpha r}}{(r+1)!r!\Gamma(\alpha r+1)}$
$=(u^{2}t^{\alpha})^{-1}\sum_{r=1}^{\infty}\frac{(u^{2}t^{\alpha})^{r}}{r!\Gamma(\alpha r+1-\alpha)},$
where we have used (2.3) at the penultimate step. This proves the first part. Relation (4.3) implies that the $l$th moment of $Z(L^{\alpha}(t))$ equals (4.2) and the second part follows. ∎
4.1 Patterned random matrices with randomly stopped walks
Let $X_{1},X_{2},\dots$ be iid rvs that are independent of the fPP $\{N^{\alpha}(t)\}_{t\geq 0}$. Consider the scaled continuous time random walk $\{c^{-\alpha/2}\sum_{k=1}^{N^{\alpha}(ct)}X_{k}\}_{t\geq 0}$, $c>0$. It is known that if $X_{k}$’s belong to the domain of attraction of standard Gaussian distribution then $\{c^{-\alpha/2}\sum_{k=1}^{N^{\alpha}(ct)}X_{k}\}_{t\geq 0}$ weakly converges to a time-changed process $\{W(L^{\alpha}(t))\}_{t\geq 0}$ in an appropriate Skorokhod topology, as $c\rightarrow\infty$ (see [23]) where $\{W(t)\}_{t\geq 0}$ is an SBM that is independent of an inverse $\alpha$-stable subordinator $\{L^{\alpha}(t)\}_{t\geq 0}$ as in Section 2.4.
In this section, we study the convergence of random matrices whose entries are randomly stopped independent random walks. Let $\{X_{i,j,k},\ i\geq 1,\ j\geq 1,\ k\geq 1\}$ be independent rvs that are independent of $\{N^{\alpha}(t)\}_{t\geq 0}$. We consider a bi-sequence $\{\mathcal{S}_{i,j,n}^{\alpha}(t),\ i\geq 1,\ j\geq 1\}_{t\geq 0}$, $n\in\mathbb{N}$ of continuous time random walks defined as follows:
$\mathcal{S}_{i,j,n}^{\alpha}(t)\coloneqq\sum_{k=1}^{N^{\alpha}(nt)}X_{i,j,k},\ t\geq 0,\ i\geq 1,\ j\geq 1,\ n\geq 1.$ (4.4)
Note that (4.4) is a fractional compound Poisson process, and is a compound Poisson process for $\alpha=1$. Let us consider the patterned random matrices $\{A_{n}^{\alpha}(t)\}_{t\geq 0}$ defined as
$A_{n}^{\alpha}(t)\coloneqq n^{-\alpha/2}(\mathcal{S}^{\alpha}_{L(i,j),n}(t))_{1\leq i,j\leq n},\ 0<\alpha\leq 1,\ n\geq 1.$ (4.5)
We also assume that $\{X_{i,j,k},\ 1\leq i\leq j,\ k\geq 1\}$ satisfies Assumption I and the link function $L$ of $A_{n}^{\alpha}(t)$, $0<\alpha\leq 1$ is symmetric, and satisfies Property B. Here also if $L$ takes its values in $\mathbb{N}$ then with abuse of notation the input sequence $\{X_{i,j,k},\ 1\leq i\leq j,\ k\geq 1\}$ is a bi-sequence $\{X_{i,k},\ i\geq 1,\ k\geq 1\}$.
In commutative probability, a Markov process, time-changed by an inverse stable subordinator, inherently produces a heavy-tailed distribution. For example, the fPP, see Section 2.4. Expecting something similar, Figures 1 and 2 illustrate the ESDs of $n^{-1/2}A_{n}^{\alpha}(t)$ for different values of $\alpha\in(0,1]$, corresponding to Wigner and Symmetric Circulant type link functions, respectively. It is observed that, for smaller values of $\alpha$, the eigenvalue spectrum is more widely spread, whereas increasing $\alpha$ leads to a contraction in the range of the $x$-axis. That is, the tails of the ESDs diminish as $\alpha$ increases.
For $\alpha=1$, entries of $A^{\alpha}_{n}(t)|_{\alpha=1}$ are $n^{-1/2}S_{L(i,j),n}(t)=n^{-1/2}\sum_{k=1}^{N(nt)}X_{L(i,j),k}$, where $\{N(t)\}_{t\geq 0}$ is a Poisson process. For each $1\leq i,j\leq n$, the increments in $\mathbb{T}_{p}(S_{L(i,j),n}(t))$ involve $X_{i,j,k}$’s indexed by Poisson points in disjoint intervals. Hence increments in $\mathbb{T}_{p}(S_{L(i,j),n}(t))$, as well as in $\mathbb{T}_{p}(A_{n}^{1}(t))$ are mutually independent. The following result on $\mathbb{T}_{p}(n^{-1/2}A_{n}^{1}(t))$ is similar to Theorem 3.3:
Theorem 4.1.
Let $A^{1}_{n}(t)=A^{\alpha}_{n}(t)|_{\alpha=1}$ be as in (4.5). For $1\leq l<\infty$, let $g(\mathbb{T}_{p}(n^{-1/2}A_{n}^{1}(t)))\coloneqq A_{r_{1},n}^{1}\dots A_{r_{l},n}^{1}$ be a monomial in indeterminates $A_{r,n}^{1}\in\mathbb{T}_{p}(n^{-1/2}A_{n}^{1}(t))$, $r=1,2,\dots,p$, $p\in\mathbb{N}$. If the limits in (3.7) exist then
$\lim_{n\rightarrow\infty}\frac{\mathbb{E}\mathrm{Trace}(A_{r_{1},n}^{1}\dots A_{r_{l},n}^{1})}{n}=\begin{cases}0,\ l=2m-1,\\ \displaystyle\prod_{\begin{subarray}{c}j=1\\ r_{j}\neq r_{i},\,j\neq i\end{subarray}}^{m}(t_{r_{j}}-t_{r_{j}-1})\sum_{w\in\mathcal{P}_{2}(2m,g)}\lim_{n\rightarrow\infty}\frac{\#\Pi(\chi_{g}(w))}{n^{1+m}},\ l=2m\end{cases}$ (4.6)
for all $m\geq 1$.
Proof.
For $0\leq s<t<\infty$, using the stationary increments property of the Poisson process,
$\mathbb{E}|S_{L(i,j),n}(t)-S_{L(i,j),n}(s)|^{l}$ $=\sum_{h\leq r=0}^{\infty}\mathbb{E}\Big|\sum_{k=h}^{r}X_{L(i,j),k}\Big|^{l}\mathbb{P}\{N(ns)=h,\,N(nt)=r\}$
$\leq\sum_{h\leq r=0}^{\infty}K_{l}(r-h)^{[l/2]}\mathbb{P}\{N(ns)=h,\,N(nt)=r\}$
$=K_{l}\sum_{r^{\prime}=0}^{[l/2]}S([l/2],r^{\prime})(n(t-s))^{r^{\prime}}\leq K_{l}^{\prime}n^{l/2}(t-s)^{l/2}$
for every $l\geq 3$ and for all $1\leq i,j\leq n$, and $\mathbb{E}(S_{L(i,j),n}(t)-S_{L(i,j),n}(s))^{2}=n(t-s)$. Here, we have used the $l$th moment of Poisson process, that is, $\mathbb{E}(N(t))^{l}=\sum_{r=0}^{l}S(l,r)t^{r}$, to get the last step, and where $S(l,r)$ is the Stirling number of second kind. Now, the proof follows similar lines to that of Theorem 3.3 (i). ∎
Remark 4.1.
For $\alpha=1$, $S_{i,j,n}(t)=\sum_{k=1}^{N(nt)}X_{i,j,k}$, where $\{N(t)\}_{t\geq 0}$ is a homogeneous Poisson process independent of $\{X_{i,j,k}\}$, and $A^{1}_{n}(t)=n^{-1/2}(S_{L(i,j),n}(t))_{1\leq i,j\leq n}$. From Theorem 4.1, for $p=1$ and $t_{0}=0$, it follows by arguments similar to those in the proof of Theorem 3.2, that the ESD of $n^{-1/2}A^{1}_{n}(t)$ converges weakly a.s. to a probability distribution whose $l$th moment is given by
$\lim_{n\rightarrow\infty}\frac{\mathbb{E}\mathrm{Trace}(A_{n}^{1}(t))^{l}}{n^{1+l/2}}=\begin{cases}0,\ l=2m-1,\ m\geq 1,\vskip 5.69046pt\\ \displaystyle t^{m}\sum_{w\in\mathcal{P}_{2}(2m)}\lim_{n\rightarrow\infty}\frac{\#\Pi(w)}{n^{1+m}},\ l=2m,\ m\geq 1.\end{cases}$ (4.7)
Thus, from (3.5), the LSD of $n^{-1/2}A^{1}_{n}(t)$ and of $n^{-1}A_{n}(t)$ (as in Theorem 3.1) are identical.
Figure 1: ESD of $n^{-1/2}A_{n}^{\alpha}(t)$ for Wigner type link function, $n=1000$, $t=1$, and with standard Gaussian step variables $X_{i,j,k}$’s.
Figure 2: ESD of $n^{-1/2}A_{n}^{\alpha}(t)$ for Symmetric Circulant type link function, $n=1500$, $t=1$, and with standard Gaussian step variables $X_{i,j,k}$’s.
Remark 4.2.
If the limits (3.7) exist for every monomial of even order then $\{n^{-1/2}A_{n}^{1}(t)\}_{t\geq 0}$ associated to $(\mathcal{M}_{n}(\mathbb{C}),\ \tau_{n}=n^{-1}\mathbb{E}\mathrm{Trace})$ algebraically converges to $\{A^{1}(t)\}_{t\geq 0}$ associated to $(\tilde{\mathcal{A}}^{1},\tilde{\tau}^{1})$, where algebra $\tilde{\mathcal{A}}^{1}$ is generated by the arbitrary union $\cup_{p\in\mathbb{N}}\cup_{\mathbb{T}_{p}}\mathcal{A}^{1}(\mathbb{T}_{p})$ with $\mathcal{A}^{1}(\mathbb{T}_{p})$ being polynomial algebra over $\mathbb{C}$, generated by the collection $\{A^{1}(t):t\in\mathbb{T}_{p}\}$. The state $\tilde{\tau}^{1}$ on it is defined similarly to (3.9). Indeed, $\tilde{\tau}^{1}(g(\{A^{1}(t):t\in\mathbb{T}_{l}\}))$ is a finite sum of limits of type (4.6) and (4.7).
If $L(i,j)=(\max\{i,j\},\min\{i,j\})$, then the limiting variables $\{A^{1}(t)\}_{t\geq 0}$ form an FBM. If L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$, then the finite dimensional moments of the limit variables agree with those of an SBM.
Remark 4.3.
Let $N(n)$ be a Poisson rv with mean $n$. Let $\{A_{k,n}\}_{k\geq 1}$ be a sequence of $n\times n$ independent random matrices defined as $A_{k,n}=n^{-1/2}(X_{L(i,j),k})_{1\leq i,j\leq n}$ for all $k\geq 1$, where the input sequence $X_{i,j,k}$’s are as in Theorem 3.1. Also, let $\{A_{k,n}\}_{k\geq 1}$ be independent of $N(n)$, and have a common symmetric link function, satisfying Property B. If the limits (3.4) exist, then the LSD of $n^{-1/2}(A_{1,n}+\dots+A_{N(n),n})$ coincides with that of $A_{k,n}$ for each $k\geq 1$. It has a probability law whose $l$th moment equals (4.7). Thus, from Remark 3.1, it follows that $n^{-1/2}(A_{1,n}+\dots+A_{N(n),n})$ and $n^{-1/2}(A_{1,n}+\dots+A_{n,n})$ have a common LSD.
Let $\{X_{i,j,k},1\leq i\leq j,\,k\geq 1\}$ be as in Theorem 3.1. Let us consider $S_{i,j,n}(t)=n^{-1/2}\sum_{k=1}^{N(nt)}X_{i,j,k}$ for $1\leq i,j\leq n$ and $t\geq 0$, and a random matrix $A_{n}(t)=(S_{L(i,j),n}(t))$. Then, it can be easily established that Theorem 4.1 holds true for the matrix valued process $\{A_{n^{\alpha}}(t)\}_{t\geq 0}$, $\alpha\in(0,1]$ for an appropriate link function $L$. We now use this fact and the next result to establish the finite dimensional algebraic convergence of $\{A^{\alpha}_{n}(t)\}_{t\geq 0}$ to the time-changed FBM $\{B^{\alpha}(t)\}_{t\geq 0}$ as defined in (4.1).
Lemma 4.1.
Let $\{N(t)\}_{t\geq 0}$ and $\{N^{\alpha}(t)\}_{t\geq 0}$ be the homogeneous and fractional Poisson processes, respectively. Let $\{X_{k}\}_{k\geq 1}$ be a sequence of independent rvs that is independent of $\{N(t)\}_{t\geq 0}$ and $\{N^{\alpha}(t)\}_{t\geq 0}$. Also, let $\{L^{\alpha}(t)\}_{t\geq 0}$ be an inverse $\alpha$-stable subordinator that is independent of both $\{N(t)\}_{t\geq 0}$ and $\{X_{k}\}_{k\geq 1}$. If $S_{n}(t)=n^{-1/2}\sum_{k=1}^{N(nt)}X_{k}$ and $S_{n}^{\alpha}(t)=n^{-\alpha/2}\sum_{k=1}^{N^{\alpha}(nt)}X_{k}$ for all $t\geq 0$, then $S^{\alpha}_{n}(t)\overset{d}{=}S_{n^{\alpha}}(L^{\alpha}(t))$, where $\overset{d}{=}$ denotes the equality in distribution.
Proof.
From [24], it follows that the inverse $\alpha$-stable subordinator is a self-similar process of index $\alpha\in(0,1)$, that is, $L^{\alpha}(ct)\overset{d}{=}c^{\alpha}L^{\alpha}(t)$ for $c>0$, and from (2.4), we have $N^{\alpha}(nt)\overset{d}{=}N(L^{\alpha}(nt))$. Therefore, $N^{\alpha}(nt)\overset{d}{=}N(n^{\alpha}L^{\alpha}(t))$ for all $t\geq 0$. This completes the proof. ∎
Theorem 4.2.
Let $\{A^{\alpha}_{n}(t)\}_{t\geq 0}$ be a matrix valued process where $A^{\alpha}_{n}(t)$ is defined in (4.5). For $p\geq 1$, $\mathbb{T}_{p}(n^{-1/2}A^{\alpha}(t))$ be the collection of increments of $\{A^{\alpha}_{n}(t)\}_{t\geq 0}$, as defined in (3.6). Let us consider a monomial $g(\mathbb{T}_{p}(n^{-1/2}A^{\alpha}(t)))=A^{\alpha}_{r_{1},n}\dots A^{\alpha}_{r_{l},n}$, where $A_{r_{j},n}^{\alpha}\in\mathbb{T}_{p}(n^{-1/2}A^{\alpha}(t))$ for each $j=1,2,\dots,l\geq 1$. Also, let $F^{\alpha}_{t_{1},\ldots t_{p}}(x_{1},\ldots,x_{p})$ be the finite dimensional distribution of an inverse $\alpha$-stable subordinator $\{L^{\alpha}(t)\}_{t\geq 0}$, that is, $F^{\alpha}_{t_{1},\ldots t_{p}}(x_{1},\ldots,x_{p})=\mathbb{P}\{L^{\alpha}(t_{1})\leq x_{1},\dots,L^{\alpha}(t_{p})\leq x_{p}\}$, $x_{1},\dots,x_{p}\in[0,\infty)$. If the limits in (3.7) exist, then
$\lim_{n\rightarrow\infty}$ $\frac{\mathbb{E}\mathrm{Trace}(A_{r_{1},n}^{\alpha}\dots A_{r_{l},n}^{\alpha})}{n}$
$=\begin{cases}0,\ l=2m-1,\ m\geq 1,\\ \displaystyle\iint_{[0,\infty)^{2m}}\prod_{\begin{subarray}{c}j=1\\ r_{j}\neq r_{i},\,j\neq i\end{subarray}}^{m}(x_{r_{j}}-x_{r_{j}-1})\sum_{w\in\mathcal{P}_{2}(2m,g)}\lim_{n\rightarrow\infty}\frac{\#\Pi(\chi_{g}(w))}{n^{1+m}}\\ \hskip 28.45274pt\cdot\mathrm{d}F^{\alpha}_{t_{r_{1}-1},t_{r_{1}},\dots,t_{r_{m}-1},t_{r_{m}}}(x_{r_{1}-1},x_{r_{1}},\dots,x_{r_{m}-1},x_{r_{m}}),\ l=2m,\,m\geq 1.\end{cases}$ (4.8)
Proof.
For a circuit $\pi$ of length $l\geq 1$, set
$S_{\pi,n}^{\alpha}=\mathbb{E}(S_{L(\pi(0),\pi(1)),n}^{\alpha}(t_{r_{1}})-S_{L(\pi(0),\pi(1)),n}^{\alpha}(t_{r_{1}-1}))\dots(S_{L(\pi(l-1),\pi(l)),n}^{\alpha}(t_{r_{l}})-S_{L(\pi(l-1),\pi(l)),n}^{\alpha}(t_{r_{l}-1})).$
By abuse of notation, let $\{r_{1},\dots,r_{l^{\prime}}\}$, $1\leq 1\leq l$ be the collection of all distinct $r_{j}$’s, and let $\{S_{i,j,n}(t)=n^{-1/2}\sum_{k=1}^{N(nt)}X_{i,j,k}\}_{t\geq 0}$ be independent of $\{L^{\alpha}(t)\}_{t\geq 0}$. Then, from Lemma 4.1, we have
$S_{\pi,n}^{\alpha}=\iint_{\mathbb{R}^{2l^{\prime}}_{+}}S_{\pi,n^{\alpha}}\mathrm{d}F^{\alpha}_{t_{r_{1}-1},t_{r_{1}},\dots,t_{r_{l^{\prime}}-1},t_{r_{l^{\prime}}}}(x_{r_{1}-1},x_{r_{1}}\dots,x_{r_{l^{\prime}}-1},x_{r_{l^{\prime}}}),$
where
$S_{\pi,n^{\alpha}}=\mathbb{E}(S_{L(\pi(0),\pi(1)),n^{\alpha}}(x_{r_{1}})-S_{L(\pi(0),\pi(1)),n^{\alpha}}(x_{r_{1}-1}))\dots(S_{L(\pi(l-1),\pi(l)),n^{\alpha}}(x_{r_{l}})-S_{L(\pi(l-1),\pi(l)),n^{\alpha}}(x_{r_{l}-1})).$
As the inverse stable subordinator is an increasing process (see [24]), for distinct indexes $r_{j}$’s, the rvs $(S_{L(\pi(0),\pi(1)),n^{\alpha}}(x_{r_{1}})-S_{L(\pi(0),\pi(1)),n^{\alpha}}(x_{r_{1}-1})),\dots,(S_{L(\pi(l-1),\pi(l)),n^{\alpha}}(x_{r_{l}})-S_{L(\pi(l-1),\pi(l)),n^{\alpha}}(x_{r_{l}-1}))$ are mutually independent due to the independent increments property of the process $\{S_{i,j,n^{\alpha}}(t)=n^{-\alpha/2}\sum_{k=1}^{N(n^{\alpha}t)}X_{i,j,k}\}_{t\geq 0}$. Thus,
$\lim_{n\rightarrow\infty}\frac{\mathbb{E}\mathrm{Trace}(A_{r_{1},n}^{\alpha}\dots A_{r_{l},n}^{\alpha})}{n}$ $=\lim_{n\longrightarrow\infty}\frac{1}{n^{1+l/2}}\sum_{\pi:\pi\text{is a circuit of length}\ l}\iint_{\mathbb{R}^{2l^{\prime}}_{+}}S_{\pi,n^{\alpha}}$
$\hskip 14.22636pt\cdot\mathrm{d}F^{\alpha}_{t_{r_{1}-1},t_{r_{1}},\dots,t_{r_{l^{\prime}}-1},t_{r_{l^{\prime}}}}(x_{r_{1}-1},x_{r_{1}}\dots,x_{r_{l^{\prime}}-1},x_{r_{l^{\prime}}}).$ (4.9)
Note that for $l=2m$ in (4.9), we have $l^{\prime}=m$. Similar to Theorem 4.1, it can be shown that $\sum_{\pi:\pi\text{is a circuit of length}\ l}\lim_{n\longrightarrow\infty}{S_{\pi,n^{\alpha}}}/{n^{1+l/2}}$ equals (4.6), which on using in (4.9), we get the required result. Also, the interchange of integral and limit in (4.9) is justified due to the dominated convergence theorem because for each $\pi$, $S_{\pi,n^{\alpha}}/n^{1+(1+\alpha)l/2}$ can be bounded by an appropriate function of $x_{r_{1}-1},x_{r_{1}},\dots,x_{r_{l}-1},x_{r_{l}}$, and all the finite dimensional joint moments of an inverse stable subordinator exist (see [31]). ∎
Remark 4.4.
For $p=1$, $t_{0}=0$, from (4.2), we get $\lim_{n\rightarrow\infty}n^{-1}{\mathbb{E}\mathrm{Trace}(A_{n}^{\alpha}(t)^{l}}=0$ whenever $l$ is odd. For $m\geq 1$, it reduces to
$\lim_{n\rightarrow\infty}\frac{\mathbb{E}\mathrm{Trace}(A_{n}^{\alpha}(t)^{2m}}{n}$ $=\int_{0}^{\infty}x^{m}\sum_{w\in\mathcal{P}_{2}(2m,g)}\lim_{n\rightarrow\infty}\frac{\#\Pi(\chi_{g}(w))}{n^{1+m}}\mathbb{P}\{L^{\alpha}(t)\in\mathrm{d}x\}$
$=\frac{t^{m\alpha}\Gamma(m+1)}{\Gamma(m\alpha+1)}\sum_{w\in\mathcal{P}_{2}(2m,g)}\lim_{n\rightarrow\infty}\frac{\#\Pi(\chi_{g}(w))}{n^{1+m}},$ (4.10)
where we have used (2.3). Thus, the EESD of $n^{-1/2}A_{n}^{\alpha}(t)$ weakly converges to a probability law whose all odd moments are zero and $2m$th moment equals (4.10). Moreover, following the proof of Theorem 3.2, it can be easily shown that the ESD of $n^{-1/2}A^{\alpha}_{n}(t)$ converges weakly a.s. to a probability law with zero odd moments and even moments are given by (4.10). In particular, the self-adjoint element $n^{-1/2}A_{n}^{\alpha}(t)$ of $(\mathcal{M}_{n}(\mathbb{C}),n^{-1}\mathbb{E}\mathrm{Trace})$ converges to an element $A^{\alpha}(t)$ of an algebra generated by $A^{\alpha}(t)$ with a state defined by $\tau((A^{\alpha}(t))^{l})=0$ for all odd $l$ and $\tau((A^{\alpha}(t))^{2m})$ equals (4.10) for each $m\geq 1$.
Remark 4.5.
Let $\{A_{k,n}\}_{k\geq 1}$ be a sequence of random matrices as defined in Remark 4.3, and it is independent of the fractional Poisson rv $N^{\alpha}(n)$. If the limits (3.4) exist, then the ESD of $A_{k,n}$ converges weakly a.s. to a probability distribution whose $l$th moment is given by (4.7) evaluated at $t=1$. Moreover, the scaled random sum $n^{-\alpha/2}(A_{1,n}+\dots+A_{N^{\alpha}(n),n})$ has an LSD whose $l$th moment is given by (4.10) at $t=1$.
The following result follows from Remark 4.4 and the results of [10]:
Corollary 4.1.
For $t\geq 0$, let $A^{\alpha}_{n}(t)$ be the random matrix as defined in (4.5). Then,
(i) for $L(i,j)=(\max\{i,j\},\min\{i,j\})$, the matrix $n^{-1/2}A^{\alpha}_{n}(t)$ in $(\mathcal{M}_{n}(\mathbb{C}),\tau_{n}=n^{-1}\mathbb{E}\mathrm{Trace})$ converges to $B^{\alpha}(t)$ (for definition see (4.1)) in polynomial algebra $\mathbb{C}\langle B^{\alpha}(t)\rangle$ with state defined as $\tau((B^{\alpha}(t))^{l})=\lim_{n\rightarrow\infty}\tau_{n}((A^{\alpha}_{n}(t))^{l})$ for all $l\geq 1$;
(ii) If L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j|| for $1\leq i,j\leq n$ then $n^{-1/2}A^{\alpha}_{n}(t)$ algebraically converges to a variable whose moments determine the probability law same as that of $W(L^{\alpha}(t))$, where $\{W(t)\}_{t\geq 0}$ is an SBM that is independent of an inverse $\alpha$-stable subordinator $\{L^{\alpha}(t)\}_{t\geq 0}$.
4.2 Symmetric circulant matrix and time-changed SBM
For $\alpha\in(0,1]$, let $\lambda_{1,n}^{\alpha}(t),\dots,\lambda_{n,n}^{\alpha}(t)$ be the eigenvalues of the random matrix $A_{n}^{\alpha}(t)$, as defined in (4.5) with link function L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$. Let $U_{n}$ be a uniform rv on $\{1,\dots,n\}$. We consider a random process $\{Y_{n}^{\alpha}(t)\}_{t\geq 0}$ defined as
$Y_{n}^{\alpha}(t)\coloneqq\lambda_{U_{n},n}^{\alpha}(t),\ \text{for all $t\geq 0$}$ (4.11)
In particular, we have
$\lambda_{r,n}^{\alpha}(t)=n^{-\alpha/2}S_{0,n}^{\alpha}(t)+2n^{-\alpha/2}\sum_{j=1}^{[n/2]}S_{j,n}^{\alpha}(t)\cos\bigg(\frac{2\pi jr}{n}\bigg),\ 0<\alpha\leq 1,\ r=1,2,\dots,n,$
where $S^{\alpha}_{j,n}(t)$ is as defined in (4.5). For $\alpha=1$, it reduces to
$\lambda_{r,n}^{1}(t)=n^{-1/2}S_{0,n}(t)+2n^{-1/2}\sum_{j=1}^{[n/2]}S_{j,n}(t)\cos\bigg(\frac{2\pi jr}{n}\bigg),\ r=1,2,\dots,n,$
where $S_{j,n}(t)=\sum_{k=1}^{N(nt)}X_{j,k}$ is as defined in Remark 4.1, and $\{N(t)\}_{t\geq 0}$ is a Poisson process with parameter one and it is independent of $X_{j,k}$’s. Let $\{L^{\alpha}(t)\}_{t\geq 0}$ be an inverse $\alpha$-stable subordinator (see Section 2.4) that is independent of $\{N(t)\}_{t\geq 0}$ and $X_{j,k}$’s. From Lemma 4.1, we get $\lambda^{\alpha}_{r,n}(t)\overset{d}{=}\lambda_{r,n^{\alpha}}^{1}(L^{\alpha}(t))$ for all $t\geq 0$. Thus,
$Y^{\alpha}_{n}(t)\overset{d}{=}Y^{1}_{n^{\alpha}}(L^{\alpha}(t)),\ t\geq 0.$ (4.12)
Theorem 4.3.
Let $Y^{\alpha}_{n}(t)$ be as defined in (4.11). Then, the process $\{n^{-1/2}Y_{n}^{\alpha}(t)\}_{t\in[0,1]}$ weakly converges to $\{W(L^{\alpha}(t))\}_{t\in[0,1]}$, where $\{W(t)\}_{t\geq 0}$ is the SBM that is independent of an inverse $\alpha$-stable subordinator $\{L^{\alpha}(t)\}_{t\geq 0}$.
Proof.
In view of Remark 4.2, following the proof of Theorem 3.4, it is easy to show that $\{n^{-1/2}Y^{1}_{n^{\alpha}}(t)\}_{t\in[0,1]}$ weakly converges to the SBM $\{W(t)\}_{t\in[0,1]}$. So, $\{(n^{-1/2}Y^{1}_{n^{\alpha}}(t),L^{\alpha}(t))\}_{t\in[0,1]}$ jointly weakly converges to $\{(W(t),L^{\alpha}(t))\}_{t\in[0,1]}$. Thus, from (4.12) and continuous mapping theorem, $\{n^{-1/2}Y_{n}^{\alpha}(t)\}_{t\in[0,1]}$ converges to $\{W(L^{\alpha}(t))\}_{t\in[0,1]}$ in an appropriate Skorokhod topology (for details see [23], Section 4.4). This completes the proof. ∎